Data for Hermann Corporation are shown below: Selling price Variable expenses Contribution margin Per Unit 140 91 $ 49 Percent of Sales 100% 65 35% Fixed expenses are $88,000 per month and the company is selling 3,000 units per month Exercise 6-5 Part 1 Required: 1-a. How much will net operating income increase (decrease) per month if the monthly advertising budget increases by $9,300 and monthly sales increase by $21,500? 1-b. Should the advertising budget be increased? Complete this question by entering your answers in the tabs below Req 1A Req 1B How much will net operating income increase (decrease) per month if the monthly advertising budget increases by $9,300 and monthly sales increase by $21,500? (Do not round intermediate calculations.) Net operating income by
Answer #1
Current Proposed
Sales (3000*140)=420,000 (420,000+21500)=441,500
Variable expenses (3000*91)=273000 (441500*65%)=$286975
Contribution margin 147,000 154,525
Fixed expenses 88000 (88000+9300)=97300
Net operating income 59000 57225

Hence Net operating income decreases by=(59000-57225)=$1775

Hene advertising budget should not be increased.